0.7 10.3 1 million G 1 million 10.2 Assume that each of these projects i s just as risky as the firms existing ets and th Which set of projects should be COST OF COMMON EQUITY The future earnings, dividends, and stock price of Carpetto Technologies Inc. are expected to grow 7% per year. Carpettos common stock currently sells for $23.00 per share, its last dividend was $2.00; and it will pay a $2.14 dividend at the end of the current year a. Using the DCF approach, what is its cost of common equity? b. at the firm may accept all the projects or only some of them. accepted? Explain. 11-6 commorn If the firms beta is 1.6, the risk-free rate is 9%, and the average return on the market is 13%, what will be the firms cost of common equity using the CAPM approach? If the firms bonds earn a return of 12%, based on the bond-yield- plus-risk-premium approach, what will be r,? Use the midpoint of the risk premium range discussed in Section 11-5 in your calculations. c. d. If you have equal confidence in the inputs used for the three phht i yte fCpt t n equity?
Answer #1
a) Cost of equity = D1/P0+g, where D1 = the next expected
dividend, g = growth rate in dividends and P0 = Current
price of the stock.
Cost of equity = 2.14/23+0.07 = 16.30%
b) Cost of equity per CAPM = risk free rate+beta*(expected market return-risk free rate).
= 9%+1.6*(13%-9%) = 15.40%
c) Mid-point range details required
Cost of equity =
d) It is the simple average of the above three.